The roots of the equation ycube+ysquare-y-1=0 are?
Answer by lwsshak3(11628)

The roots of the equation ycube+ysquare-y-1=0 are?
y^3+y^2-y-1=0
y^2(y+1)-(y+1)=0
(y+1)(y^2-1)=0
(y+1)(y+1)(y-1)=0
roots are:
x=1
and
x=-1(multiplicity 2)